OpenStudy (anonymous):

If you were an Algebra 1 instructor and were creating a test on factoring trinomials of the form ax2 + bx + c, what do you think would be the easiest way to create a trinomial that can be factored? Provide one unique example.

OpenStudy (amistre64):

x^2 + 0x + 0

OpenStudy (amistre64):

just a hunch

OpenStudy (anonymous):

I used 4 and 9 so 4+9=13 and 4*9=36 6*6=36 the equations should be... 6x^2+4x+9x+6 6x^2+13x+6 right?